OpenStudy (anonymous):

In one data set, the mean is smaller than the median. What is the symmetry of the data?

OpenStudy (anonymous):

The data is skewed to the left. The data is symmetric. It is not possible to determine the symmetry of the data. The data is skewed to the right.

OpenStudy (dan815):

data is skewed to the right

OpenStudy (anonymous):

can you explain why please?

OpenStudy (dan815):

suppose there are your values 1 2 81 82 83 84 85 86 <--- median being the middle number is 83 However the mean due to the 1,2 will decrease the value as a whole 1+2+81+82...86 -------------- =584/8=73 8

OpenStudy (dan815):

give me a second to think though, if we can find a counter example where, even though the data is skewed to the right, and the mean is above the median. Then we will have to conclude it is not possible to determine the symmetry of the data.

OpenStudy (dan815):

okay here I found a site that explaining it well

OpenStudy (dan815):

http://www.mathsisfun.com/data/skewness.html

OpenStudy (dan815):

sorry I misunderstood the meaning of "skew", it should be skewed to the left, when the values of data are more on the right side.. go figure

OpenStudy (dan815):

So again, the answer was "Skewed to the left" not "right".